OpenStudy (nuccioreggie):

Arlene is showing her work in simplifying −7.9 + 8.2 − 3.4 + 2.1. Identify any errors in her work or in her reasoning. Write feedback to Arlene explaining the mistake she made and how to correct her work. Step Work Reasoning −7.9 + 8.2 − 3.4 + 2.1 Original problem 1 −7.9 + 8.2 + (−3.4) + 2.1 Additive inverse 2 −7.9 + (−3.4) + 8.2 + 2.1 Associative property 3 (−7.9 + (−3.4)) + (8.2 + 2.1) Commutative property 4 −11.3 + 10.3 Simplify 5 −1 Simplify

OpenStudy (nuccioreggie):

@DanJS

OpenStudy (nuccioreggie):

@iambatman

OpenStudy (nuccioreggie):

@Ahsome ? anything

OpenStudy (ahsome):

There is no errors

OpenStudy (ahsome):

are*

OpenStudy (ahsome):

Keep doing each step manually, and see where it slips up: \[−7.9 + 8.2 − 3.4 + 2.1=-1\] \[−7.9 + 8.2 + (−3.4) + 2.1=-1\] \[−7.9 + (−3.4) + 8.2 + 2.1=-1\] \[(−7.9 + (−3.4)) + (8.2 + 2.1)=-1\] \[−11.3 + 10.3 =-1\]